Monopolistically competitive market with N firms

Consider a monopolistically competitive market with N firms. Each firm's business opportunities are described by the following equations:
Demand: Q=100/N-P
Marginal Revenue: MR=100/N-2Q
Total cost: TC=50+Q(squared)
Marginal Cost: MC=2Q

a. How does N, the number of firms in the market, affect each firms demand curve? Why.
b. How many units does each firm produce? (The answer to this and the next two questions depend on N.)
c. What price does each firm charge?
d. How much profit does each firm make?
e. In the long run, how many firms will exist in this market?
Answer #1

(a) Each firms demand curve is given by the following equation: 100 The demand functionisithal nterceptofthe demand u 100 ;(c) Substitute each firms production amount in individual firms demand function. 25 100 100 25 100-25 75 S75 Therefore, eac(e) In the long run, firms make zero profit. 1250 72-50-0 1250 25 N°-25 Therefore, only five fims will exist in the long run.
Answer #2

a) With increase in number of firms(N) , the quantity produced decreases as N has inverse relation with Q as seen in the demand equation.

b) For monopolistic competition, each firm produces at a point where ,

MR = MC

100/N - 2Q = 2Q

or, 100/N = 4Q

or, Q* = 25/N

c) We know,

Q = 100/N - P

or, 25/N = 100/N - P

or, 1/N (25-100) = -P

or, 75/N = P*

d) Total Profit = TR - TC = 100Q/N - 4 Q2

= 100(25)/N2 - 4(125/N2)

= 2000/N2    

e)

MAXIMIZE Profit = d(TR)dQ - d(TC)/dQ = 0

or,100/N - 4Q - 4Q

or, 100Q/N - 8Q = 0

or, 100Q/N = 8Q

or, N * = 100/8 = 12.5 i.e 12